answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

What is the probability that certain genotypes and phenotypes will occur?

This depends entirely on the genotype of the parents. The probability of getting a specific genotype is the probability of getting the correct allele from mother (1/2) multiplied by the probability of getting the correct allele from father (1/2) multiplied by the number of ways this can occur. The probability of getting a phenotype, if the phenotype is dominant, is the sum of the probability of getting two dominant alleles, and the probability of getting one dominant allele. If the phenotype is recessive, the probability is equal to the probability of getting two recessive alleles.

How many 3 digit numbers can be formed by 3 4 5 6 7 8 without repetition?

120.

There are 6 digits. If we pick the digits in order, there are

6 possible digits for the first digit

5 remaining digits for the second digit

4 remaining digits for the third digit.

6*5*4 = 120.

What is the probability that a couple with 3 children will have at least 1 girl?

The easiest way of calculating this is to find the probability that all three are boys, as this is the only arrangement that does not fit the criteria. Then work out the answer by taking this away from 1. Probability that all three are boys = 1/2 x 1/2 x 1/2 = 1/8. probability of there being at least one girl is 1 - 1/8 = 7/8 or 87.5%

Why do you need to calculate standard deviation and relative error?

to ensure your experiment is precise and to prevent error to happen during experiment

Are complementary events also mutually exclusive events?

Yes, they are. Mutually exclusive events cannot occur together. Complementary events cannot occur together either because an event and its complement are the negative of each other.

When a single card is drawn from an ordinary 52-card deck find the probability of getting a diamond?

There are 13 diamonds in a 52-card deck

so the probability to get a diamond is 13/52=1/4=0.25

What is the probability of drawing a black card from a 52 card deck and flipping heads on a coin?

Half the cards in a standard pack are black.

Therefore the probability of drawing a black card is 1/2.

Half the sides of a coin are "heads" so again the probability is 1/2.

Therefore the probability you will both draw a black card and flip heads = 1/2 * 1/2 = 1/4.

What is the probability of being dealt 3 cards and all of them being diamonds?

There are 13 diamonds. Three cards are dealt. The probability of all of them being diamond is (13/52)(12/51)(11/50) = 1716/132600 = 11/850

Is -0.0031 probability?

No, a probability is a number between 0 (included) and 1 (included)

What is the probability it will rain our of four days?

The probability is 1.

It is always raining - any time of the day or night, any day of the year - somewhere on Earth.

What is required in a binomial distribution?

A number of trials, each of which has only two outcomes: these are usually termed "success" and "failure". The trials must be independent and the probability of success must remain constant.

What is the probability of spinning an odd number 8 1 7 2 6 5 4 3?

8 numbers in total, 4 of those numbers are odd (1, 7, 5 and 3)

so 4/8

= 1/2

aka 50%, 0.5

What is the probability of tomorrow being a Sunday?

There are 7 days in a week. One of those is Sunday. Therefore the probability at any given time that tomorrow will be Sunday is 1/7.

What does even distribution mean?

It means that the probability density function is symmetric about 0.

Why will a density curve always give a probability of zero?

It will not. For the interval (x, x+dx) it may well give a non-zero probability.

With a continuous distribution, the probability of any particular value is always 0. What the probability density function gives is the probability that the variable is NEAR the selected value.

Uses of t-score and z-score in the classroom?

Z scores are used for standardized testing done by most school districts.

It is the most common way of standardizing data. IQ scores can be standardized using z scores. The mean is 100 and the standard deviation is 15.

You use the t score when the sample is small, <30 often. Many behavior ratings use t scores.

Why are the events of spinning a spinner and rolling a number cube independent events?

Because the number cube is not sentient enough to know the result of the spinner and modify its own outcome accordingly. And conversely, the outcome of the spinner is not affected by the roll of the cube.

5000 students participated in a certain test yielding a result that follows the normal distribution with means of 65 points and standard deviation of 10 points.?

5000 students participated in a certain test yielding a result that follows the normal distribution with means of 65 points and standard deviation of 10 points.

(1) Find the probability of a certain student marking more than 75 points and less than 85 points inclusive.

(2) A student needs more than what point to be positioned within top 5% of the participants in this test?

(3) A student with more than what point can be positioned within top 100 students?

i dont understand the question.. could you help me??pls....